# theorem]Construction

### v1

@@ 4. $H$-percolating graphs
We also looked at the threshold $p_{c}(n,K_{k})$ for when a random graph $G(n,p)$ is $K_{k}$-percolating. Above this threshold, we saw that typical edges of $K_{n}$ play the role of the final edge $e_{d}$ of some $F_{d}$ generated by a $K_{k}$-chain after removing the edges $e_{i}$. Moreover, we can take that this chain is logarithmic in size and so random graphs above the threshold will also percolate fast. In fact, this was studied in detail by Gunderson, Koch and Przykucki [51] who showed that just above the threshold $p_{c}(n,K_{k})$ the random graph will a.a.s. $K_{k}$-percolate in $O(\log\log n)$ steps and this time decreases as $p$ increases.
@@
Note that every $(2,1)$-inseparable graph is $3$-connected as it has at least five vertices and cannot be disconnected by deleting two vertices. Furthermore every $4$-connected graph is $(2,1)$-inseparable. Indeed, if the removal of an edge $e$ and two vertices $z_{1},z_{2}$ disconnected a $4$-connected graph $H$ then so would the removal of $z_{1},z_{2}$ and an appropriate endpoint of $e$ (as $H$ has at least five vertices). The notion of $(2,1)$-inseparability plays a crucial role with respect to chain constructions. Indeed, it was precisely this property of $K_{5}$ that was used to prove Observation 2.4 showing that simple chains do not induce unwanted copies of $K_{5}^{-}$.
@@ 4.2. $H$-percolating random graphs
For general $H$ with $\delta(H)\geq 2$ and $v(H)\geq 4$, Bartha and Kolesnik [17] proved a lower bound $n^{-1/\lambda_{*}(H)+o(1)}$ which is governed by a more subtle function $\lambda_{*}(H)=\min\frac{e_{H}-e_{F}-1}{v_{H}-v_{F}}$, where the minimum is taken over all subgraphs $F\subset H$ with $2\leq v_{F}<v_{H}$. They show that for balanced graphs $\lambda^{*}=\lambda$, so together with the upper bound of [12] the asymptotics of the exponent of the threshold is determined for all balanced graphs. Recently Bartha, Kronenberg, and Kolesnik [16] proved $H=G(k,1/2)$ is balanced a.a.s. and hence the critical exponent for the $H$-process percolating is equal to $1/\lambda(H)$ for “most” infection rules $H$.
@@ 4.3. Slow $H$-percolating graphs
Let $u\in U$, and let $w_{1},\ldots,w_{\delta(H)-1}\in W$ be pairwise distinct $G$-neighbours of $u$. For $w\in W\setminus\{w_{1},\ldots,w_{\delta(H)-1}\}$, any map $\varphi:V(H)\to V(G)$ that sends a vertex of minimum degree to $u$, its neighbourhood to $\{w,w_{1},\ldots,w_{\delta(H)-1}\}$, and the remaining vertices of $H$ to the rest of $W$ is an embedding of $H$ minus an edge into $\langle G\rangle_{H}$. Thus $uw$ completes a copy of $H$ and there must be a complete bipartite graph between $U$ and $W$ in $\langle G\rangle_{H}$. All edges in $\binom{U}{2}$ then complete some copy of $H$ with other vertices chosen from $W$ and so indeed $U\cup W$ is a clique in $\langle G\rangle_{H}$. ∎
@@
This puts us in good stead but we still need to find a way to attach the set $U$ for each new gadget to $\delta(H)-1$ vertices in $V(G)$ without interfering with the process on $G$. To do this, we simply add an independent set $I$ to $G$ and consider the auxiliary $H$-percolating graph $G^{\prime}$ with linearly many edges as a graph on vertex set $V(G)\cup I$. Therefore our gadgets will force that $V(G)\cup I$ hosts a complete graph and we can attach the vertex sets $U$ to vertices in $I$ to trigger the percolation on the individual gadgets. One more idea is needed to guarantee that the edges between $U$ and $I$ for different gadgets do not interact with each other to give unwanted copies of $H$ minus an edge. We will use a bipartite graph of large girth to place these edges between gadgets and $I$, recalling that the girth of a graph is the length of the shortest cycle in the graph. Specifically, we need the following.
@@
Bipartite graphs with many edges and high girth have been studied thoroughly in the literature, see for example [63, 64, 50]. In fact, we will use them again in our construction of line chains in Section 7.4. We note that for our purposes we do not require the full strength of those results because the parameters $k$ and $d$ are fixed constants. Indeed, Lemma 4.8 can be obtained easily by considering a random bipartite graph with parts $X$ of size $n$ and $Y^{\prime}$ of size $(2k+1)n$, with each edge present independently with probability $p=n^{-1+1/(2k)}$. By an application of Chernoff’s theorem, a.a.s. all vertices in $Y^{\prime}$ will have degree at least $d$ and by Markov’s inequality, a.a.s. there will be at most $n$ cycles of length at most $k$. Therefore a.a.s. both these events happen and we can obtain the graph $B$ in Lemma 4.8 by first deleting from $Y^{\prime}$ any vertex that participates in a cycle of length at most $k$ and then deleting further vertices of $Y^{\prime}$ and edges to get $Y$ of size exactly $kn$ with each vertex incident to $d$ edges. Finally then, we give the details of the proof of Theorem 4.2
@@
Let $H$ be a $(2,1)$-inseparable $k$-vertex graph and $G$ an $n$-vertex graph with running time $\tau_{H}(G)=M_{H}(n)$. We construct a super-graph $\tilde{G}\supseteq G$ such that $\tau_{H}(\tilde{G})\geq\tau_{H}(G)$ and $\tilde{G}$ is $H$-percolating. Introduce an independent set $I$ of $n$ new vertices, and let $G^{\prime}$ be an auxiliary graph with vertex set $V(G)\cup I$ such that $|E(G^{\prime})|\leq 2kn$ and $\langle G^{\prime}\rangle_{H}=K_{2n}$. Let $\Gamma$ be the graph from Lemma 4.7 with $\ell:=v(\Gamma)$ vertices, and for each $f\in G^{*}:=G^{\prime}\setminus\langle G\rangle_{H}$ introduce a set $W^{f}$ of $\ell-4$ new vertices. Fix an edge $e_{\tau}\in G_{\tau}\setminus G_{\tau-1}$ and place a copy $\Gamma^{f}$ of $\Gamma$ on $W^{f}\cup e_{\tau}\cup f$ such that for some subset $U^{f}\subseteq W^{f}$, (1)–(3) of Lemma 4.7 hold for $e_{\tau},f,U^{f}$ (with $e_{\tau}$ playing the role of $e$).
Consider an auxiliary bipartite graph $B$ between $I$ and $\{U^{f}:f\in G^{*}\}$ such that
@@
Let $F$ be a copy of $H$ minus an edge in $\tilde{G}_{t-1}$. Suppose first that $V(F)$ intersects $I$. As $N_{\tilde{G}_{t-1}}(v)=N_{\tilde{G}}(v)\subseteq\cup_{f\in G^{*}}U^{f}$, the connectedness of $F$ and the inequality $\operatorname{dist}_{\Gamma^{f}}(e_{\tau}\cup f,U^{f})\geq k$ imply that $V(F)\subseteq I\cup\left(\cup_{f\in G^{*}}W^{f}\right)$. Every $v\in V(F)\cap I$ has at most one $F$-neighbour in $U^{f}$ for each $f\in G^{*}$, and hence $|\{f\in G^{*}:U^{f}\cap N_{F}(v)\neq\varnothing\}|\geq\delta(F)>2$, using that $H$ is $(2,1)$-inseparable in the last inequality here. Moreover, for every $f\in G^{*}$ such that $U^{f}\cap V(F)\neq\varnothing$, we must have that there are at least two $B$-neighbours of $U^{f}$ in $I$ as otherwise we could disconnect $F$ by removing one vertex, contradicting the $(2,1)$-inseparability of $H$. Therefore we have that $B[(I\cap V(F))\cup\{U^{f}:U^{f}\cap V(F)\neq\varnothing\}]$ has minimum degree greater than two and therefore contains a cycle. Each vertex in this cycle corresponds to (at least one) vertex of $V(F)$ and these vertices are distinct. Hence the cycle has length at most $v(F)=k$ which contradicts the assumption that the girth of $B$ is at least $k+1$.
@@ 5.2. Random graphs
Theorem 5.1 already establishes that cliques of size at least $5$ are far from unique in having maximum running time that is (almost) quadratic. Our next result shows that in fact almost all infection rules on $n$ vertices have quadratic running time. In fact, unlike Theorem 5.1, where the density is a large constant, for typical infection rules $H$, a very sparse density turns out to be sufficient to cause quadratic behaviour.
@@
It is natural to ask whether this threshold is sharper than presented here. The 0-statement, asserting that $M_{H}(n)$ is constant when $p$ is small, follows from the fact that an isolated edge occurs a.a.s. (see for example [30, Theorem 5.4]). The $1$-statement, giving quadratic running time for $p$ much larger than $\log k/k$, is more delicate and in particular, we use that $H=G(k,p)$ is self-stable, i.e. the $H$-process is stabilises on $H$ as the starting graph. For this we apply a result of Kim, Sudakov and Vu [58]. Whilst it is believable that results on the asymmetry of $G(k,p)$ can be strengthened to give that $H=G(k,p)$ is a.a.s. self-stable as soon as $p>\log k/k$, we did not find such a result in the literature and it is not immediate how one could prove such a statement. Nonetheless, we believe it is likely that our proof can be pushed to give $M_{H}(n)=\Omega(n^{2})$ a.a.s. already when $p>c_{1}\log k/k$ for any $c_{1}>1$. The range $\log k/2k<p<\log k/k$ seems to be the most interesting and it is unclear what to expect for the behaviour of $M_{H}(n)$ with $H=G(k,p)$.
@@
With the additional condition $p<\frac{\log k}{k}$, the infection rule $H$ in the problem will a.a.s. have one giant component and some isolated vertices, which can be ignored from the perspective of running times. There will also be vertices of degree one which suggests that our chain constructions will be useless for proving lower bounds on $M_{H}(n)$ as copies of $H$ minus an edge will be abundant and so no analogue of the key property 1 from Lemma 2.3 can hold. Conversely, as we will see in Section 7.1, the existence of degree 1 vertices alone is not enough to be able to prove effective upper bounds on running times.
@@ 5.3.1. The square of the Hamilton cycle
The first result is about the $4$-regular family of the squares of Hamilton cycles. To construct the square of the Hamilton cycle one orders the $k$ vertices cyclically and connects any two of them whose distance is at most two.
@@ 5.3.2. The wheel
Our second result establishes an almost quadratic lower bound for the maximum running time of a sequence of graphs with average degree strictly less than four. The wheel graph $W_{k}$ is defined to be the graph with $k+1$ vertices obtained by taking a cycle of length $k$ and adding a new vertex that is adjacent to all vertices of the cycle. The striking feature of this example is that we can also couple the lower bound with a non-trivial upper bound. In particular this introduces a completely new behaviour of possible maximum running times, which is sub-quadratic but larger than any polynomial of degree smaller than $2$. This can be seen as evidence for the validity of Conjecture 2.5 which predicts a similar behaviour for $K_{5}$.
@@ 5.4. Proof technique: Chain constructions
Avoiding local copies $F$ of HeH-e for some edge $e\in H$ essentially boils down to showing that we cannot place such a copy $F$ on the edges of a simple $H$-chain $(H_{i},e_{i})_{i\in[\tau_{0}]}$ without placing it on $V(H_{i})$ for some $i\in[\tau_{0}]$. For $H=K_{5}$, we noticed this via Observation 2.4. For general infection rules $H$, there is no guarantee that we cannot separate a copy $F$ of HeH-e across a vertex cut of size $2$ (given by some edge $e_{i}$ of the chain). This motivates the definition of $(2,1)$-inseparability from Section 4, which provides a simple condition on $H$ that suffices to avoid the local unwanted copies of HeH-e.
As with $K_{5}$, in many cases we can forbid the presence of crossing unwanted copies $F$ of HeH-e by placing dilation chains $\mathcal{H}$ according to an appropriate dilation set $A$ which avoids solutions to certain linear equations and placing one simple $H$-chain corresponding to each dilation. Through Fact 3.5, we showed that an unwanted copy of $K_{5}^{-}$ leads to a triangle spanning more than one chain. In some of our applications in this section, for example for the wheel graph $W_{k}$, we cannot guarantee this. We can however work with a slightly weaker concept.
@@
The non-monochromatic cycles from Definition 5.7 can still be avoided by taking a dilation set $A$ which is of Behrend type, as in Theorem 3.2, and choosing the constant bounding the size of the coefficients of the equations to be sufficiently large in terms of $k=v(H)$. We therefore arrive at the following theorem.

### v2

@@ 4. $H$-percolating graphs
We also looked at the threshold $p_{c}(n,K_{k})$ for when a random graph $G(n,p)$ is $K_{k}$-percolating. Above this threshold, we saw that typical edges of $K_{n}$ play the role of the final edge $e_{d}$ of some $F_{d}$ generated by a $K_{k}$-chain after removing the edges $e_{i}$. Moreover, we can take that this chain is logarithmic in size and so random graphs above the threshold will also percolate fast. In fact, this was studied in detail by Gunderson, Koch and Przykucki [51] who showed that just above the threshold $p_{c}(n,K_{k})$ the random graph will $a.$s. $K_{k}$-percolate in $O(\log\log n)$ steps and this time decreases as $p$ increases.
@@
Note that every $(2,1)$-inseparable graph is 33-connected as it has at least five vertices and cannot be disconnected by deleting two vertices. Furthermore every 44-connected graph is $(2,1)$-inseparable. Indeed, if the removal of an edge $e$ and two vertices $z_{1},z_{2}$ disconnected a 44-connected graph $H$ then so would the removal of $z_{1},z_{2}$ and an appropriate endpoint of $e$ (as $H$ has at least five vertices). The notion of $(2,1)$-inseparability plays a crucial role with respect to chain constructions. Indeed, it was precisely this property of $K_{5}$ that was used to prove Observation 2.4 showing that simple chains do not induce unwanted copies of $K_{5}^{-}$.
@@ 4.2. $H$-percolating random graphs
For general $H$ with $\delta(H)\geq 2$ and $v(H)\geq 4$, Bartha and Kolesnik [17] proved a lower bound $n^{-1/\lambda_{*}(H)+o(1)}$ which is governed by a more subtle function $\lambda_{*}(H)=\min\frac{e_{H}-e_{F}-1}{v_{H}-v_{F}}$, where the minimum is taken over all subgraphs $F\subset H$ with $2\leq v_{F}<v_{H}$. They show that for balanced graphs $\lambda^{*}=\lambda$, so together with the upper bound of [12] the asymptotics of the exponent of the threshold is determined for all balanced graphs. Recently Bartha, Kronenberg, and Kolesnik [16] proved $H=G(k,1/2)$ is balanced $a.$s. and hence the critical exponent for the $H$-process percolating is equal to $1/\lambda(H)$ for “most” infection rules $H$.
@@ 4.3. Slow $H$-percolating graphs
Let $u\in U$, and let $w_{1},\ldots,w_{\delta(H)-1}\in W$ be pairwise distinct $G$-neighbours of $u$. For $w\in W\setminus\{w_{1},\ldots,w_{\delta(H)-1}\}$, any map $\varphi:V(H)\to V(G)$ that sends a vertex of minimum degree to $u$, its neighbourhood to $\{w,w_{1},\ldots,w_{\delta(H)-1}\}$, and the remaining vertices of $H$ to the rest of $W$ is an embedding of $H$ minus an edge into $\langle G\rangle_{H}$. Thus uwuw completes a copy of $H$ and there must be a complete bipartite graph between $U$ and $W$ in $\langle G\rangle_{H}$. All edges in $\binom{U}{2}$ then complete some copy of $H$ with other vertices chosen from $W$ and so indeed $U\cup W$ is a clique in $\langle G\rangle_{H}$. ∎
@@
This puts us in good stead but we still need to find a way to attach the set $U$ for each new gadget to $\delta(H)-1$ vertices in $V(G)$ without interfering with the process on $G$. To do this, we simply add an independent set II to $G$ and consider the auxiliary $H$-percolating graph $G^{\prime}$ with linearly many edges as a graph on vertex set $V(G)\cup I$. Therefore our gadgets will force that $V(G)\cup I$ hosts a complete graph and we can attach the vertex sets $U$ to vertices in II to trigger the percolation on the individual gadgets. One more idea is needed to guarantee that the edges between $U$ and II for different gadgets do not interact with each other to give unwanted copies of $H$ minus an edge. We will use a bipartite graph of large girth to place these edges between gadgets and II, recalling that the girth of a graph is the length of the shortest cycle in the graph. Specifically, we need the following.
@@
Bipartite graphs with many edges and high girth have been studied thoroughly in the literature, see for example [63, 64, 50]. In fact, we will use them again in our construction of line chains in Section 7.4. We note that for our purposes we do not require the full strength of those results because the parameters $k$ and $d$ are fixed constants. Indeed, Lemma 4.8 can be obtained easily by considering a random bipartite graph with parts $X$ of size $n$ and $Y^{\prime}$ of size $(2k+1)n$, with each edge present independently with probability $p=n^{-1+1/(2k)}$. By an application of Chernoff’s theorem, $a.$s. all vertices in $Y^{\prime}$ will have degree at least $d$ and by Markov’s inequality, $a.$s. there will be at most $n$ cycles of length at most $k$. Therefore $a.$s. both these events happen and we can obtain the graph $B$ in Lemma 4.8 by first deleting from $Y^{\prime}$ any vertex that participates in a cycle of length at most $k$ and then deleting further vertices of $Y^{\prime}$ and edges to get $Y$ of size exactly knkn with each vertex incident to $d$ edges. Finally then, we give the details of the proof of Theorem 4.2
@@
Let $H$ be a $(2,1)$-inseparable $k$-vertex graph and $G$ an $n$-vertex graph with running time $\tau_{H}(G)=M_{H}(n)$. We construct a super-graph $\tilde{G}\supseteq G$ such that $\tau_{H}(\tilde{G})\geq\tau_{H}(G)$ and $\tilde{G}$ is $H$-percolating. Introduce an independent set II of $n$ new vertices, and let $G^{\prime}$ be an auxiliary graph with vertex set $V(G)\cup I$ such that $|E(G^{\prime})|\leq 2kn$ and $\langle G^{\prime}\rangle_{H}=K_{2n}$. Let $\Gamma$ be the graph from Lemma 4.7 with $\ell:=v(\Gamma)$ vertices, and for each $f\in G^{*}:=G^{\prime}\setminus\langle G\rangle_{H}$ introduce a set $W^{f}$ of $\ell-4$ new vertices. Fix an edge $e_{\tau}\in G_{\tau}\setminus G_{\tau-1}$ and place a copy $\Gamma^{f}$ of $\Gamma$ on $W^{f}\cup e_{\tau}\cup f$ such that for some subset $U^{f}\subseteq W^{f}$, (1)–(3) of Lemma 4.7 hold for $e_{\tau},f,U^{f}$ (with $e_{\tau}$ playing the role of $e$).
Consider an auxiliary bipartite graph $B$ between II and $\{U^{f}:f\in G^{*}\}$ such that
@@
Let $F$ be a copy of $H$ minus an edge in $\tilde{G}_{t-1}$. Suppose first that $V(F)$ intersects II. As $N_{\tilde{G}_{t-1}}(v)=N_{\tilde{G}}(v)\subseteq\cup_{f\in G^{*}}U^{f}$, the connectedness of $F$ and the inequality $\operatorname{dist}_{\Gamma^{f}}(e_{\tau}\cup f,U^{f})\geq k$ imply that $V(F)\subseteq I\cup\left(\cup_{f\in G^{*}}W^{f}\right)$. Every $v\in V(F)\cap I$ has at most one $F$-neighbour in $U^{f}$ for each $f\in G^{*}$, and hence $|\{f\in G^{*}:U^{f}\cap N_{F}(v)\neq\varnothing\}|\geq\delta(F)>2$, using that $H$ is $(2,1)$-inseparable in the last inequality here. Moreover, for every $f\in G^{*}$ such that $U^{f}\cap V(F)\neq\varnothing$, we must have that there are at least two $B$-neighbours of $U^{f}$ in II as otherwise we could disconnect $F$ by removing one vertex, contradicting the $(2,1)$-inseparability of $H$. Therefore we have that $B[(I\cap V(F))\cup\{U^{f}:U^{f}\cap V(F)\neq\varnothing\}]$ has minimum degree greater than two and therefore contains a cycle. Each vertex in this cycle corresponds to (at least one) vertex of $V(F)$ and these vertices are distinct. Hence the cycle has length at most $v(F)=k$ which contradicts the assumption that the girth of $B$ is at least $k+1$.
@@ 5.2. Random graphs
Theorem 5.1 already establishes that cliques of size at least 55 are far from unique in having maximum running time that is (almost) quadratic. Our next result shows that in fact almost all infection rules on $n$ vertices have quadratic running time. In fact, unlike Theorem 5.1, where the density is a large constant, for typical infection rules $H$, a very sparse density turns out to be sufficient to cause quadratic behaviour.
@@
It is natural to ask whether this threshold is sharper than presented here. The 0-statement, asserting that $M_{H}(n)$ is constant when $p$ is small, follows from the fact that an isolated edge occurs $a.$s. (see for example [30, Theorem 5.4]). The 11-statement, giving quadratic running time for $p$ much larger than $\log k/k$, is more delicate and in particular, we use that $H=G(k,p)$ is self-stable, i.e. the $H$-process is stabilises on $H$ as the starting graph. For this we apply a result of Kim, Sudakov and Vu [58]. Whilst it is believable that results on the asymmetry of $G(k,p)$ can be strengthened to give that $H=G(k,p)$ is $a.$s. self-stable as soon as $p>\log k/k$, we did not find such a result in the literature and it is not immediate how one could prove such a statement. Nonetheless, we believe it is likely that our proof can be pushed to give $M_{H}(n)=\Omega(n^{2})$ $a.$s. already when $p>c_{1}\log k/k$ for any $c_{1}>1$. The range $\log k/2k<p<\log k/k$ seems to be the most interesting and it is unclear what to expect for the behaviour of $M_{H}(n)$ with $H=G(k,p)$.
@@
With the additional condition $p<\frac{\log k}{k}$, the infection rule $H$ in the problem will $a.$s. have one giant component and some isolated vertices, which can be ignored from the perspective of running times. There will also be vertices of degree one which suggests that our chain constructions will be useless for proving lower bounds on $M_{H}(n)$ as copies of $H$ minus an edge will be abundant and so no analogue of the key property 1 from Lemma 2.3 can hold. Conversely, as we will see in Section 7.1, the existence of degree 1 vertices alone is not enough to be able to prove effective upper bounds on running times.
@@ 5.3.1. The square of the Hamilton cycle
The first result is about the 44-regular family of the squares of Hamilton cycles. To construct the square of the Hamilton cycle one orders the $k$ vertices cyclically and connects any two of them whose distance is at most two.
@@ 5.3.2. The wheel
Our second result establishes an almost quadratic lower bound for the maximum running time of a sequence of graphs with average degree strictly less than four. The wheel graph $W_{k}$ is defined to be the graph with $k+1$ vertices obtained by taking a cycle of length $k$ and adding a new vertex that is adjacent to all vertices of the cycle. The striking feature of this example is that we can also couple the lower bound with a non-trivial upper bound. In particular this introduces a completely new behaviour of possible maximum running times, which is sub-quadratic but larger than any polynomial of degree smaller than 22. This can be seen as evidence for the validity of Conjecture 2.5 which predicts a similar behaviour for $K_{5}$.
@@ 5.4. Proof technique: Chain constructions
Avoiding local copies $F$ of HeH-e for some edge $e\in H$ essentially boils down to showing that we cannot place such a copy $F$ on the edges of a simple $H$-chain $(H_{i},e_{i})_{i\in[\tau_{0}]}$ without placing it on $V(H_{i})$ for some $i\in[\tau_{0}]$. For $H=K_{5}$, we noticed this via Observation 2.4. For general infection rules $H$, there is no guarantee that we cannot separate a copy $F$ of HeH-e across a vertex cut of size 22 (given by some edge $e_{i}$ of the chain). This motivates the definition of $(2,1)$-inseparability from Section 4, which provides a simple condition on $H$ that suffices to avoid the local unwanted copies of HeH-e.
As with $K_{5}$, in many cases we can forbid the presence of crossing unwanted copies $F$ of HeH-e by placing dilation chains $\mathcal{H}$ according to an appropriate dilation set AA which avoids solutions to certain linear equations and placing one simple $H$-chain corresponding to each dilation. Through Fact 3.5, we showed that an unwanted copy of $K_{5}^{-}$ leads to a triangle spanning more than one chain. In some of our applications in this section, for example for the wheel graph $W_{k}$, we cannot guarantee this. We can however work with a slightly weaker concept.
@@
The non-monochromatic cycles from Definition 5.7 can still be avoided by taking a dilation set AA which is of Behrend type, as in Theorem 3.2, and choosing the constant bounding the size of the coefficients of the equations to be sufficiently large in terms of $k=v(H)$. We therefore arrive at the following theorem.